# Neurosymbolic Association Rule Mining from Tabular Data

### v3

@@ 1 Introduction
To address this research problem, we make the following contributions: i) a novel neurosymbolic ARM method - Aerial+ (Section 3); ii) two comprehensive evaluations of Aerial+ (Section 4) on 5 real-world tabular datasets (Kelly et al., 2023) that demonstrate Aerial+’s superiority over seven baselines both in knowledge discovery and downstream classification tasks.
@@ 2 Related Work
Interpretable ML. Besides knowledge discovery, ARM is widely used in rule-based interpretable ML models, which is the standard approach to high-stake decision-making (Rudin, 2019). Examples include associative classifiers such as CBA (Liu et al., 1998), rule sets, and rule list learners (Letham et al., 2015; Angelino et al., 2018) that construct rule-based classifiers from a set of pre-mined rules or frequent itemsets via ARM and class labels. Since these models rely on ARM to pre-mine rules or frequent itemsets, the excessive number of rules and long execution times carry over to downstream interpretable ML tasks, further increasing computational costs. All these methods work with exhaustive ARM approaches such as FP-Growth to pre-mine frequent itemsets and rules. Numerous versions of FP-Growth have also been proposed to alleviate the aforementioned issues such as Guided FP-Growth (Shabtay et al., 2021) for ARM with item constraints, parallel FP-Growth (Li et al., 2008) and FP-Growth on GPU (Jiang and Meng, 2017) for better execution times.
@@ 3.2 Autoencoder Architecture and Training Stage
Let $F=\{f_{1},f_{2},...,f_{k}\}$ be a set of $k$ features in a tabular dataset (e.g. columns in a table) and $f_{i}^{1...c_{i}}$ represent categories (possible values) for feature $f_{i}$ $(1\leq i\leq k)$ where $c_{i}=|f_{i}|$ indicates the number of categories for feature $f_{i}$. To represent such a tabular dataset as an ARM problem as defined in Section 2, we define the full set of items II as consisting of all possible categories of all features $I=\{f_{1}^{1},...,f_{1}^{c_{1}},...,f_{k}^{1},...,f_{k}^{c_{k}}\}$. Each transaction $T$ in the tabular dataset (e.g., a row in a table) is represented as an itemset $T\subset I$, where each feature contributes exactly one item: $\forall i\in\{1,...,k\},\exists!j\in\{1,...,c_{i}\}\text{ such that }f_{i}^{j}\in T$. Using one-hot encoded representation, $f_{i}^{1...c_{i}}=\{0,1\}$, with 0 and 1 indicating the absence or the presence of a feature category in a given transaction.
@@ 3.3 Rule Extraction Stage
Intuition. Autoencoders can learn a neural representation of data and this representation includes the associations between the feature categories. We hypothesize that the reconstruction ability of autoencoders can be used to extract these associations. After training, if a forward run on the trained model with a set of marked categories AA results in successful reconstruction (high probability) of categories $C$, we say that marked features AA imply the successfully reconstructed features $C$, such that $A\rightarrow{C}\setminus A$ (no self-implication).
@@ 8       $\mathbf{p}\leftarrow AE(\mathbf{v})$;
Algorithm. The rule extraction process is given in Algorithm 1. Line 1 stores input feature categories into $\mathcal{F}$. Lines 2-14 iterate over the number of antecedents $i$ and line 3 generates an $i$-feature combination $\mathcal{F}$. Lines 4-13 iterate over the feature combinations $\mathcal{C}$. For each combination $S$, line 5 creates a vector with uniform probabilities per feature category, e.g., $[0.5,0.5,0.33,0.33,0.33]$ vector in the example given above. Line 6 creates a set of test vectors where a combination of feature categories in $S$ are marked per test vector. This corresponds to the $[1,0,0.33,0.33,0.33]$ vector in the example where $weather(warm)$ was marked. Lines 7-13 iterate over the test vectors. Line 8 performs a forward run on the trained autoencoder AEAE with each test vector $v$. Lines 9-11 compare the output probabilities corresponding to $S$ with a given antecedent similarity threshold $\tau_{a}$, and the algorithm continues with high probability $S$ values. Lines 12-13 compare the output probabilities for the categories in $\mathcal{F}$ that are not in $S$ already ($F\setminus S$), with a given consequent similarity threshold $\tau_{c}$ and stores categories with high probability. Finally, line 14 removes the low support categories from $\mathcal{F}$ so they are ignored in line 3 of the next iteration.
@@ Appendix A Runtime Complexity Analysis of Aerial+
    Performs a forward pass on the trained autoencoder AEAE in $O(|\mathcal{F}|)$.
@@ F.4 Rule Extraction Stage
Forward pass and output interpretation. Next, we feed the input AA through the trained autoencoder to obtain output:
@@ F.5 Interpretation of the formal justification
  • Probing with test vectors. Rule extraction relies on probing the decoder with a constructed input AA that sets $x_{i}=c$ deterministically and others to uniform noise. The resulting output $\bar{x}$ estimates $\mathbb{P}(x_{j}=c^{\prime}\mid x_{i}=c)$ (Equation 6), enabling the discovery of association rules.

### v5

@@ 1 Introduction
To address this, we make the following contributions: i) a novel neurosymbolic ARM method - Aerial+ (Section 3); ii) two comprehensive evaluations of Aerial+ (Section 4) on 5 real-world tabular datasets (Kelly et al., 2023) that demonstrate Aerial+’s superiority over seven baselines in knowledge discovery and downstream classification tasks.
@@ 2 Related Work
Interpretable ML. Besides knowledge discovery, ARM is widely used in rule-based interpretable ML models, which is the standard approach to high-stake decision-making (Rudin, 2019). Examples include associative classifiers such as CBA (Liu et al., 1998), rule sets, and rule list learners (Letham et al., 2015; Angelino et al., 2018) that construct rule-based classifiers from a set of pre-mined rules or frequent itemsets via ARM and class labels. Since these models rely on ARM to pre-mine rules or frequent itemsets, the excessive number of rules and long execution times carry over to downstream interpretable ML tasks and increases computational costs. All these methods work with exhaustive ARM approaches such as FP-Growth to pre-mine frequent itemsets and rules. Numerous versions of FP-Growth have also been proposed to alleviate these issues such as Guided FP-Growth (Shabtay et al., 2021) for ARM with item constraints, parallel FP-Growth (Li et al., 2008) and FP-Growth on GPU (Jiang and Meng, 2017) for better execution times.
@@ 3.2 Autoencoder Architecture and Training Stage
Let $F=\{f_{1},f_{2},...,f_{k}\}$ be a set of $k$ features in a tabular dataset (e.g. columns in a table) and $f_{i}^{1...c_{i}}$ represent categories (possible values) for feature $f_{i}$ $(1\leq i\leq k)$ where $c_{i}=|f_{i}|$ indicates the number of categories for feature $f_{i}$. To represent such a tabular dataset as an ARM problem as defined in Section 2, we define the full set of items $I$ as consisting of all possible categories of all features $I=\{f_{1}^{1},...,f_{1}^{c_{1}},...,f_{k}^{1},...,f_{k}^{c_{k}}\}$. Each transaction $T$ in the tabular dataset (e.g., a row in a table) is represented as an itemset $T\subset I$, where each feature contributes exactly one item: $\forall i\in\{1,...,k\},\exists!j\in\{1,...,c_{i}\}\text{ such that }f_{i}^{j}\in T$. Using one-hot encoded representation, $f_{i}^{1...c_{i}}=\{0,1\}$, with 0 and 1 indicating the absence or the presence of a feature category in a given transaction.
@@ 3.3 Rule Extraction Stage
Intuition. Autoencoders can learn a neural representation of data and this representation includes the associations between the feature categories. We hypothesize that the reconstruction ability of autoencoders can be used to extract these associations. After training, if a forward run on the trained model with a set of marked categories $A$ results in successful reconstruction (high probability) of categories $C$, we say that marked features $A$ imply the successfully reconstructed features $C$, such that $A\rightarrow{C}\setminus A$ (no self-implication).
@@ 8       $\mathbf{p}\leftarrow AE(\mathbf{v})$;
Algorithm. The rule extraction process is given in Algorithm 1. Line 1 stores input feature categories into $\mathcal{F}$. Lines 2-14 iterate over the number of antecedents $i$ and line 3 generates an $i$-feature combination $\mathcal{F}$. Lines 4-13 iterate over the feature combinations $\mathcal{C}$. For each combination $S$, line 5 creates a vector with uniform probabilities per feature category, e.g., $[0.5,0.5,0.33,0.33,0.33]$ vector in the example given above. Line 6 creates a set of test vectors where a combination of feature categories in $S$ are marked per test vector. This corresponds to the $[1,0,0.33,0.33,0.33]$ vector in the example where $weather(warm)$ was marked. Lines 7-13 iterate over the test vectors. Line 8 performs a forward run on the trained autoencoder $AE$ with each test vector $v$. Lines 9-11 compare the output probabilities corresponding to $S$ with a given antecedent similarity threshold $\tau_{a}$, and the algorithm continues with high probability $S$ values. Lines 12-13 compare the output probabilities for the categories in $\mathcal{F}$ that are not in $S$ already ($F\setminus S$), with a given consequent similarity threshold $\tau_{c}$ and stores categories with high probability. Finally, line 14 removes the low support categories from $\mathcal{F}$ so they are ignored in line 3 of the next iteration.
@@ Appendix A Runtime Complexity Analysis of Aerial+
    Performs a forward pass on the trained autoencoder $AE$ in $O(|\mathcal{F}|)$.
@@ F.4 Rule Extraction Stage
Forward pass and output interpretation. Next, we feed the input $A$ through the trained autoencoder to obtain output:
@@ F.5 Interpretation of the formal justification
  • Probing with test vectors. Rule extraction relies on probing the decoder with a constructed input $A$ that sets $x_{i}=c$ deterministically and others to uniform noise. The resulting output $\bar{x}$ estimates $\mathbb{P}(x_{j}=c^{\prime}\mid x_{i}=c)$ (Equation 6), enabling the discovery of association rules.